Mr Shahrezza expected return from an investment are as follows: RM 80,000 per year for the first five years, RM 10,000 at the year end of 6 and RM 100,000 at the end of year 7. If the interest rate is at 8%, what is the present value of the cash flows by today?


1
Expert's answer
2020-07-15T10:03:34-0400

"NPV=\frac{80,000}{(1+0.08)^1}+\frac{80,000}{(1+0.08)^2}+\frac{80,000}{(1+0.08)^3}+\frac{80,000}{(1+0.08)^4}+\frac{80,000}{(1+0.08)^5}+\frac{10,000}{(1+0.08)^6}+\frac{100,000}{(1+0.08)^7}=384,067.54"